Find four consecutive even integers with a sum of 196.
A rectangle has width 52 units and perimeter 200 units. Find the length.

Answer by ankor@dixie-net.com(22740)

Find four consecutive even integers with a sum of 196.
n + (n+2) + (n+4) + (n+6) = 196
combine like terms, do some basic algebra
:
A rectangle has width 52 units and perimeter 200 units. Find the length.
2L + 2(52) = 200
simplify, divide by 2
L + 52 = 100
you should be able to finish this